Instrument Log    TDMA

TDMA            C1       Lamont, OK

Date

Technician

Comments

4/19/16

MG

Instrument needs to be shut down and restarted due to a scheduled Patch Release at 1300 GMT.  At 1511 GMT I shutdown and reset computer to install a Patch as per BCR-02040 RTS at 1519 GMT.

2/19/16

MG

On 2-16-2016 IT department issued a patch release which caused the instrument to shut down and not properly restart. Once I restarted the program the QH down blower failed. I took apart the Blower and tested the DC/DC converter and replaced the blower fan and reassembled. I then used silicone sealant to make the housing airtight and once the sealant was dry I reinstalled the fan. While instrument was down I filled the water container and emptied the water trap. RTS 1600 GMT

2/16/16

MG

Due to IT department computer needed rebooted for patch releases. Rebooted Computer at 16:00 RTS 16:13 Once rebooted the QH down Blower fan went out.

2/12/16

MG

One of the Blowers was making a loud screeching sound.I stopped the program at 16:32 and contacted mentor. He instructed to restart program and reinsured me that it was working properly and to let instrument run despite the sound. If blower fails then we will replace it. RTS 16:15 GMT

1/19/16

MG

IT department installed software patches on the instrument computer and it rebooted itself at 13:17 GMT but never started running the instrument program. At 2010 GMT I restarted the program and it came online and began normal operations. RTS at 2010 GMT.

1/8/16

MG

On 12-18-2015 the TDMA water supply stopped bubbling. Per Mentors request I changed out the power supply for a solenoid valve which temporally fixed the problem. later changed the valve itself. this did not solve the problem so i put the original valve back on the instrument. After extensive troubleshooting on the mentor and my part I discovered that the Water trap container had a very small crack on the top of the container. This prevented the instrument from creating a vacuum and not allowing the relative humidity to build. I retrofitted a new water trap that works great and now the TDMA is bubbling wonderfully. After all this was repaired the relative humidity still would not climb so I had to repair the water tank heater. all is well with the TDMA and RTS was 21:00 GMT on 1-8-2016

12/31/15

MG

12-27-2015 at 00:00 GMT SGP experienced a Ice storm that caused a power outage. the TDMA was drained after power went out. 12-31-2015 I refilled the water reservoir and restarted the program all is well with the TDMA RTS 21:38 GMT

12/21/15

MS

Big water reservoir was not bubbling during PM rounds.  Shut down instrument at 1614 GMT to locate the issue and did not find one.  TDMA water reservoir is still not functional.  Restarted program at 1616 GMT, RTS 1631.

12/18/15

MS

Water bottle was not bubbling proper.  Shut instrument down to investigate hoses to water bottle at 1535 GMT.  No problem was found during inspection.  RTS at 1537 GMT.

12/17/15

MS

Heater was off and water was not bubbling properly.  Shut off instrument at 1704 GMT.  Checked the hoses and the heater going into the water bottle.  Did not see any problems.  Heater came back on and water is bubbling proper now. Turned instrument back on at 1724 GMT.  RTS at 1724 GMT.

12/15/15

MG

Signal amplifier failed for proportional control valve but instrument was still operational. Refer to CM on 11/20/2015 for power outage on 11/17/2015 from 0700-1435 GMT. Replaced signal amplifier for proportional control valve. RTS 1635 GMT.

11/20/15

MG

Instrument was down.  One valve not responding and water reservoir not bubbling.  Worked with mentor over the phone to T/S problem.  Measured voltage input and output at amplifier.  0.6V and 24V measured, which is not the proper voltage for the solenoid valve.  Also disconnected the humidity inlet line to check for vacuum, which there was none.  Mentor will ship replacement valve and amplifier.  TDMA is running but not to its fullest capability.  RTS at 1630 GMT.

11/18/15

MG

A 7 hour power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/15. Instrument was down. Valve not responding and the water reservoir is not bubbling. Worked with mentor over the phone trying to diagnose the problem with the TDMA. Measured voltage output and input at amplifier .6 volts in and 24 out is not the proper voltage to the solenoid valve. I also disconnected the humidity in hose to check for a vacuum which there was not. Mentor is shipping a new valve and amplifier. TDMA is running but not to its fullest capabilities RTS 16:30

11/17/15

MG

SGP experienced a power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/15. I had to Reboot the computer at 1515 TDMA was still showing errors and not running properly mentor asked me to wait for further instruction so he could diagnose the problem.

10/23/2015

MG

Sheath Flow blower not working correctly. I shut down TDMA per mentors request until he could diagnose problem. Disconnected and reconnected operating signal wire to blower and it came back to life. Down at 2100 on 10-19-2015 RTS 10-23-2015 at1424

10/20/2015

MG

Per mentors request I shut down TDMA on 10-20-2015 at 0837 UTC. *Alicat Mass Flow meter Part #H12-5LPM-O-O Serial # 7652 was switched to a newer version, Alicat Mass Flow meter Part # M-5SLPM-D Serial #122969.  RTS at 0900 on 10-20-2015 all is well and will monitor performance. 

9/28/15

KT

The pump reading for CE 44 was reading below 10 psi and was failing. At 1300 GMT removed the old failing pump and installed a newly rebuilt pump. After powering up the new pump the CE 44 reading climbed up to 18 psi. System is online and operational. RTS at 1315 GMT.

9/17/15

PD/MG

APS has been returned from vendor after repairs and calibration and needs to be installed.  Installed *APS, powered up and turned on pumps.  TSI model 1109125, *SN 70815064, WD58069.  RTS at 1900 GMT.

9/16/15

PD/MG

Computer is up but Labview is not running.  Started Labview program.  RTS at 1450 GMT.

9/14/15

PD/MG

Notified by Mentor CPC readings are low or not registering at all.  Butanol level is good.  Drained small amount of butanol from CPC internal reservior and observed no water.  Further troubleshooting found 3 way valve between aerosol inlet line and atomizer solution system was turned to atomizer.  Mentor was unable to turn valve using software.  Removed 3 way valve and manually turned valve to correct position to allow inlet air.  Reinstalled valve and started program.  Mentor later reported CPC values are now in range.   Valve was left disconnected from controller so it would not move back to atomizer position.  RTS at 1609 GMT.

9/10/15

PD

I have been troubleshooting low Qsh_up values with Mentor for a while.  Yesterday a blockage in DMA was removed that may have been causing the issue.  Humidifier heater was turned on and system started yesterday after blockage was removed.  Today at 1427 GMT, I realized HV_up and HV_dn switches were off.  Stopped program, turned on switches and started program.  RTS at 1442 GMT.

9/9/15

PD

Qsh_up values have been very low.  Troubleshooting with Mentor to try to isolate the cause.  Mentor suspects there is a blockage of either butanol or NaCl buildup inside the DMA that is causing the restriction.  Connected tubing from CE44 vacuum line at bulkhead to one of the connectors at the bottom of the DMA.  Conducted numerous tests opening connections at different parts of the system to try to isolate the blockage.  Testing confirmed blockage was inside the DMA.  Consulted with Mentor and connected air compressor to bottom connector of DMA while leaving the other bottom and both top connectors open.  Flowed < 10 psi pressure through DMA and alternated closing two of the three open connections.  Repeated process with air connected to the other DMA bottom connector and both top.  Did not see anything come out of the DMA but blockage was removed.  Reinstalled all removed tubing lines, turned on humidifier heater and started program.  Notified Mentor.  RTS at 2119 GMT.

9/9/15

PD

Qsh_up values are low.  CE 44 vacuum line at bulkhead was directly connected to the bottom of DMA and allowed to run for 24 hours to try to get any suspected butanol inside DMA removed and dried out.  Qsh_up value was 0.021 yesterday prior to change and 0.036 after.  This morning it was .042 prior to putting system back to normal configuration and 0.32 after.  Stopped program at 1519 GMT and reconnected at tubing between bulkhead and DMA for normal operations.  RTS at 1544 GMT.

9/8/15

PD

Qsh_up value is currently 0.021 with a setpoint of 0.600.  Mentor suspects either butanol is still inside the DMA or possibly a salt buildup that is restricting flow.  Mentor request connecting CE44 vacuum line directly to connector on bottom of DMA where butanol was previously found.  Stopped program and connected CE44 vacuum line just inside bulkhead directly to DMA bottom connector.  Qsh_up value was 0.021 prior to change and 0.036 after.  Will notify Mentor and let system run for a couple of hours and check for improvement.  RTS at 1446 GMT.

9/3/15

PD

Qsh_up value remains low at .270.  Inlet line to blower was left disconnected since last week since Mentor was on site the week before to allow any Butanol in the lines to dry out.  Mentor suspects since that did not fix the issue, there may be salt deposits partially blocking inlet line.  Shut system down at 1933 GMT.  Troubleshooting found the only thing that significantly effects the Qsh_up value is when MFC inlet connector is disconnected from the DMA.  Qsh_up value then increased to .852.  Suspect blockage is inside DMA.  Emailed Mentor results of troubleshooting and consulting on how to proceed.  RTS at 2043 GMT.

8/28/15

PD

Inline static strips need to be changed.  Stopped program at 1438 GMT.  R/R static strips.  Observed evidence that salt from atomizer solution had been in the static strip enclosure.  Cleaned salt out of enclosure.  Removed nafion dryer upstream of static strips and observed salt deposits created some blockage of tubes at both ends of nafion.  Cleared blockages and reinstalled.  Qsh_up value increased from .275 to .335 after cleaning nafion.  RTS at 1603 GMT.

8/25/15

PD

Mentor is on site and R/R mass flow meter with loaner provided by Mentor upstream of CPC.  Mentor request MFC sent for repair and calibration.  Will submit PR for repair and calibration.  Alicat Scientific Model # V-5LPM-D, SN 24376.

8/13/15

PD

Instrument is not operational.  Mentor is planning a trip to SGP and wants to troubleshoot some issues to have more information on identifying all issues and have a better idea of items needed to facilitate repairs.  Worked with Mentor who Radmined into system to check voltages, flows, etc.  Mentor will write and email and develop plan to repair instrument.  RTS at 1825 GMT.

8/10/15

PD

Flow values are out of range for APS (Aerodynamic Particle Sizer).  Removed unit and troubleshooting found one of the internal pumps has failed.  After consulting with Mentor, unit will be sent to TSI for repair and calibration.  TSI model 1109125, SN 70815064, WD 58069.  RTS at 1619 GMT.

8/6/15

PD

Pumps need to be replaced and rebuilt.  R/R *CE and VE pumps.  Old CE model 2688CE44, SN 010500000564.  New CE model 2688CE44, SN 10050000844.  Old VE model 2688VE44, SN 060500002019.  New model 2688VE44, SN 040900004569.  RTS at 1417 GMT.

8/6/15

PD

CE vacuum pump line is breaking repeatedly at the pump connector.  Plastic line is brittle and needs to be replaced.  R/R tubing with Synflex.  Changed pumps while system was down.  RTS at 1417 GMT.

8/5/15

PD

CE vacuum pump inlet line broke at the pump connector.  Removed Swagelok nut and ferrules, cut line and reinstalled.  Vacuum line is getting brittle.  Will replace tomorrow with Synflex tubing.  RTS at 1449 GMT.

8/4/15

PD           

CE vacuum pump inlet line broke at the pump connector.  Removed Swagelok nut and ferrules, cut line and reinstalled.  RTS at 1510 GMT.

7/22/2015

MS

Program was not running during DSView check at 1407 GMT.  One of the pumps had a broken hose. Reset program at 1407 GMT and that did not work, so shutdown computer at 1407 GMT and brought it back up.  Started program and RTS at 1437 GMT. Also fixed a broken pump line as well.

7/21/2015

MS

Noticed during PM rounds that the program had shut off.  Restarted program at 1426 GMT. RTS at 1441 GMT.

6/18/2015

MS

Scheduled shut down on the TDMA program to clean out the heater reservoir container.  Shut down at 1429 GMT.     Restarted the TDMA program and heater and RTS at 1444 GMT.

6/17/2015

MS

Ground_TDMA.vi program has stopped at 1517 GMT.     Restarted Ground_TDMA.vi program at 1517 GMT.  RTS at 1532 GMT.

6/12/2015

MS

TDMA program had stopped running at 1349 GMT.  Restarted TDMA program at 1823 GMT.  RTS at 1838 GMT.

6/4/15

KT

Due to a scheduled power shut down the TDMA system went down. At 1741 GMT after the transformer upgrade was performed the power was restored and the instrument was brought back online and was operational.

5/21/2015

MS

Program on the TDMA had stopped running at 1402 GMT.  Restarted TDMA program and RTS at 1417 GMT.

2/18/15

MS

Program was not running at 1405 GMT.   Restarted program at 1420 GMT.

1/20/15

KT

During rounds the program was not running. At 1525 GMT rebooted the computer program. At 1534 GMT system is online and operational. RTS at 1540 GMT.

11/28/14

PD

Labview program is running but an error is shown: Error occurred at Open/Create/Replace File in Write Spreadsheet String.vi->Write to Spreadsheet File (DBL).vi->ground_tdma.vi.  NI-488: DMA hardware error detected.  Stopped/restarted LabView program.  RTS at 1447 GMT.

11/26/14

PD

Still in the process of troubleshooting CPC flow and counts issue.  Butanol is not visible in CPC window and instrument is showing zero counts.  Connected CPC butanol fill line and filled instrument with butanol.  Program was active but not running (stop button was gray).  Humidifier heater is turned off and CPC inlet line downstream of mass flow meter is disconnected exposing CPC to room air.  CPC started showing counts above 30,000 after adding butanol.  RTS at 1536 GMT.

11/20/14

PD/KT

Troubleshooting with Mentor on CPC showing zero counts.  Mentor suggest removing CPC and inspecting nozzle and critical orifice for blockage and verify optic block is not flooded with butanol.  Shut down and removed CPC.  Found nozzle blocked with salt buildup from atomizer solution.  Flushed nozzle with distilled water and cleared with canned air.  Cleaned small amount of salt from inside top of nozzle body with Q-tip and distilled water.  Installed CPC, left inlet line disconnected, filled with butanol until visible in sight glass, and started program.  CNC showed instantaneous counts of several hundred thousand.  Reconnected inlet line and flow indicator LED went out.  Disconnected inlet line again downstream of flowmeter and flow indicator came on.  Emptied atomizer reservoir and filled it with distilled water only.  Notified Mentor of what was found and current state of instrument.  T/S and consulting with Mentor at this time concerning where other blockages may be, suspected issue with CPC inlet flowmeter, and how to control amount of NaCL being ingested by system during midnight scans.  RTS at 2150 GMT.

11/14/14

PD

TDMA CPC is not showing any particle counts.  CPC butanol was drained yesterday and allowed to pull air for 24 hours in attempt to dry out instrument.  At 1536 GMT, added butanol until visible in sight glass, started program and allowed system to run until CPC temperature stabilized.  No particle counts were shown even when inlet tube was disconnected exposing instrument to sample room air.  Notified Mentor and was advised to drain butanol and allow system to run over the weekend.  Drained butanol at 2116 GMT and started program.  RTS at 2131 GMT

11/13/14

PD/KT

CPC has been showing 0 counts.  Mentor suspects optics may be flooded with butanol and requests CPC be drained and allowed to dry for 24 hours.  Shut down system at 1521 GMT.  Drained butanol and restarted system.  RTS at 1546 GMT.

11/12/14

PD/KT

CPC is showing 0 counts.  Mentor requests disconnecting inlet line exposing instrument to room air to see if counts are shown.  Disconnected inlet line at 2210 GMT.  No counts were shown.  Reconnected line and notified Mentor.

10/31/14

CAM

Instrument mentor advised that the upstream humidity (RH Up) was reading low.  Confirmed RH Up was only 77% when set value was 90%.  Turned vacuum pumps off at 1727 GMT.  Removed primary Nafion dryer and inspected for blockage.  Found partial blockage at each end of the sample line.  Used compressed air to remove blockage and reinstalled Nafion dryer.  Restored power to pumps at 1754 GMT.  RH Up now reading 90%, the same as the set point.  Notified mentor.

10/27/14

MS

Human error, closed out program by accident at 1254 GMT.  Restarted program at 1309 GMT.

10/15/14

CAM/TMG

The new computer’s BIOS needs to be edited so that the computer will power up automatically following a power outage.  Stopped the program at 2019 GMT.  Edited BIOS and restarted.  RTS at 2026 GMT.

9/25/14

CAM

The repaired original CPC was returned from vendor and needed to be installed in place of the loaner.  R/R *CPC at 1745 GMT.  Will keep loaner here on site until mentor confirms repaired unit is functioning properly.  System is ready for restart once mentor finishes reloading the software and drivers made necessary by the computer replacement earlier this week.  *Old – 70520330  *New - 36  (Md: 3762, WD54676) 

9/22/14

CAM/TMG

A new Advantech model SYS-4U-BTO computer was acquired to replace the failed TMDA computer.  Removed hard drives and PCI cards from the old computer and installed in the new.  Installed new *computer in the rack and powered on at 2040 GMT.  Mentor will reload the software and drivers remotely and then the instrument will be powered back on and returned to service.

*New – ACE8003304  (WD80882)

9/15/14

CAM

TDMA computer power supply has failed and cannot be repaired.  The power supply is no longer available for purchase so the computer must be replaced.  Removed Supermicro *computer to IT department in order to swap hard drives and cards to a replacement computer.    TDMA will remain offline pending successful configuration of new computer.

*Old – 1012007591  (WD54675)

9/11/14

CAM

The CPC failed on 9/3 and mentor sent a loaner instrument, which arrived today.  R/R *CPC with loaner instrument at 2030 GMT.  The TDMA computer would then not power on.  Suspect power supply has failed.  Removed power supply to electronics lab for troubleshooting.  System will remain offline pending computer repair.

*Old – 36  (Md: 3762, WD54676)  *New - 70520330

9/3/14

CAM

Instrument mentor called and indicated there were no counts from the CPC.  Check of instrument found laser LED was out.  Suspect the laser has failed but mentor suggested drying the instrument out over night to see if that would correct the problem.  Shut TDMA down at 1830 GMT but left pump running.  Removed inlet sample line from CPC and installed a HEPA filter open to room air.  Allowed CPC to dry overnight but there was no change in status the following day.  Mentor will send a replacement CPC, at which time the failed unit will be returned to vendor for repair.  Instrument shut down

8/25/14

KT

The pump reading for CE44 was reading 0. At 1420 GMT during inspection found the tubing that connects to the pump was cracked and repaired. Powered the pump on and the value went to 10. System is online and operational. RTS at 1437 GMT.

6/5/14

CAM

Mentor requested instrument and pumps be powered back on, but HV left off.  Powered on as requested at 1818 GMT.

6/4/15

CAM

Mentor requested instrument and pumps be powered off for the night.  Powered off at 2300 GMT.

6/4/14

CAM

Mentor requested instrument and pumps be powered back on, but HV left off.  Powered on as requested at 1800 GMT.

6/3/14

CAM

Mentor called and advised he was having issues with the HV.  Requested instrument be powered off overnight.  Powered instrument and pumps off at 2200 GMT.

6/2/14

CAM

Computer upgrade to Windows 7 was completed and computer reinstalled at 1343 GMT.  Mentor will reload all software and update drivers.

5/28/14

CAM

Computer needs to be upgraded from Windows XP to Windows 7.  Mentor shut system down remotely and computer was then removed at 1900 GMT and taken to IT lab.  Will reinstall computer when upgrade is complete.

5/27/14

CAM

Found the UPS that powers the AOS system and comms equipment in the AOS trailer was down.  TDMA has its own UPS and was still running, but may have been partially affected by the outage. 

5/19/14

CAM

Instrument mentor notified Site Ops the atomizer was not functioning properly.  Found the atomizer solution level was down just below the pickup tube.  Refilled atomizer bottle with NaCl / Distilled water solution at 1250 GMT.

5/14/14

CAM

Qsh down flow is reading much lower than the set point. T/S found both the blower and the DC-DC converter inside the blower box had failed.  Replaced blower with a new spare and replaced DC-DC converter with one from a blower assembly than had been previously removed from the Ambient State.  Sealed the blower box with silicone and allowed to dry.  Reinstalled blower assembly and RTS at 2227 GMT.  Flow rate returned to normal but the pre-RH reading was still high at 75%.  This may or may not be due to mentor shutting parts of the system off while blower assembly was removed.  Notified mentor and awaiting reply.

5/13/14

CAM

Shut the TDMA down at 1811 GMT so the batteries could be replaced in the UPS.  Restarted TDMA and RTS at 1835 GMT.

4/8/14

PAD/KT

TDMA computer needs additional memory to accept Windows7 upgrade.  It currently has two 256MB memory cards installed.  Stopped Labview program and shut down computer. Removed 256MB memory cards and installed two 1GB cards.  Powered up computer.  RTS at 1953 GMT. 

1/13/14

CAM

Atomizer solution bottle was empty.  Filled bottle with NaCl solution and RTS at 2245 GMT.

11/26/13

MS

Had to replace the loaner heater with a new heater, so the loaner could be returned.  Removed loaner heater and replaced with new heater at 1620 GMT.

11/18/13

MS

Water bottle heater needed to be replaced.  Installed spare heater until new heater arrives.  RTS at 1533 GMT

10/9/13

CAM

One of the Nafion dryer critical orifices has been getting plugged with NaCl residue lately.  Instrument mentor requested the stainless tubing connecting the critical orifice to the supply line be removed and the orifice left open to filtered cabinet air.  Removed the tubing as requested and installed a Balston filter on the orifice inlet.  RTS at 1530 GMT.

9/16/13

CAM

TDMA pre-RH is running high once again.  Removed Nafion dryer critical orifice and found it was obstructed with a small amount of NaCl residue.  Cleared critical orifice and RTS at 1620 GMT.  Pre-RH dropped from 54% to 37% within a couple of minutes and continued to drop.  RTS status O.K.

9/4/13

CAM

Instrument mentor requested the atomizer valve be fully closed and then opened two turns.  Adjusted valve as requested at 2120 GMT.  Noted that valve had previously been opened approximately 7-8 turns.

9/4/13

CAM

Instrument mentor advised that the pre-RH value was high again and also the particle counts have been excessively high for the past few days.  Instrument had already been powered down remotely by mentor.  Removed the Nafion dryer critical orifice at 1914 GMT and found it was once again clogged with a small spec of NaCl residue.  Cleaned critical orifice and reinstalled.  Powered instrument on and started program.  Pre-RH value started dropping as it should, but Qsh input was only 0.025 with a set point of 0.8.  Powered instrument back down.  Removed flow meter and found it too was clogged with NaCl residue.  Cleaned flow meter and reinstalled.  Powered instrument on at 2014 GMT.  Pre-RH started out at 38% and began dropping.  Within a few minutes it was down to normal reading of 19%.  Qsh input was also matching the set point, indicating that cleaning the flow meter corrected the problem.  RTS status O.K.

9/3/13

CAM

Received new replacement blower and need to install in the blower assembly for the ambient state.  Installed Ametek blower SN 012288 and removed failed blower SN 001903.  Need to seal the box with epoxy before it is reinstalled in system.

8/31/13 –

9/3/13

CAM

Received a phone call on Saturday (8/31) from shift operator who advised the TDMA calibration compressor was running non-stop and tank pressure was zero.  Advised operator to unplug compressor at that time and it would be looked at on 9/3 after the holiday.  Checked compressor on 9/3 and found the cam assembly that drives the piston had come loose from its mount, causing severe damage before finally falling completely out.  The connecting rod cracked completely through where it attaches to the cam.  There was also damage to the cylinder wall and motor shaft where the belt normally rides.  Removed compressor and installed a spare backup in its place at 1400 GMT on 9/3.  Will contact ANL Purchasing to discuss options for possibly getting the compressor repaired under warranty or purchasing a new compressor.

8/28/13

CAM

Instrument mentor advised the RH was running high once again.  Removed Nafion dryer critical orifice and found it was partially obstructed.  Cleared critical orifice and RTS at 1833 GMT.  RH dropped from 56% to 37% within a couple of minutes and continued to drop.  RTS status O.K.

8/15/13

CAM

RH values continue to be high, running around 45%.  Removed Nafion dryer critical orifice at 2110 GMT for inspection and found it was clogged with NaCl residue.  Cleaned critical orifice and returned to service at 2125 GMT.  RH value started dropping steadily and was down to 30% within a few minutes.  By the next day RH value had returned to normal range of around 19%.

8/14/13

CAM

Instrument mentor advised that swapping the Nafion dryer earlier resulted in slight improvement in Rh values, but further troubleshooting was necessary.  Requested vacuum pumps be swapped out.  Vacuum gauge readings were -16.5 inHg and -16 inHg prior to shutting down.  Stopped program at 2112 GMT and replaced vacuum pumps.  Restarted program and RTS at 2119 GMT.  Gauges now read -19 inHg and -17.5 inHg.  Mentor will evaluate new data for improvement and then determine if old pumps should be rebuilt or not.

8/14/13

CAM/KT

Instrument mentor advised there were issues with the Rh and requested the Nafion dryer be inspected.  Stopped the program and powered TDMA off at 1900 GMT.  Removed Nafion dryer and inspected.  Unable to see through the sample tube, but could blow compressed air through.  Removed outermost collars and found that the epoxy filler between the Nafion tubes had deteriorated and small pieces had actually fallen out.  Tube ends looked discolored and deformed, with some appearing nearly closed off.  At mentor’s request, removed the Nafion dryer from the Ambient State and installed it in the TDMA.  Put the damaged dryer in the Ambient State as a place holder to keep the lines sealed off.  Swapped the critical orifice and other fittings between the two that were necessary for the plumbing to match up.  Restored power and started program at 2035 GMT.  Notified mentor maintenance was complete.

Nafion dryer Model # PD-07018T-12SS. 

Old TDMA dryer SN: 714-0305-36  New TDMA dryer SN: 713-0705-5

7/16/13

CAM

Instrument mentor advised the counts are still a bit noisy during the midnight calibration.  Opened the atomizer valve 1 full turn as requested by mentor at 2230 GMT.

7/9/13

CAM

Instrument mentor advised the counts were still low during the midnight calibration, even after changing the atomizing solution.  Opened the atomizer valve 1 full turn as requested by mentor at 2000 GMT.

7/8/13

KT

Instrument mentor requested that Atomizer solution be removed and replaced and to turn the Atomizer valve 1 full turn clockwise.  At 1900 GMT stopped the program and removed and replaced the Atomizer solution and also turned the atomizer valve 1 full turn clockwise. RTS at 1915 GMT.

7/2/13

CAM

Mentor advised the counts were lower than normal during the midnight calibrations.  Requested atomizer valve be opened further by one full turn.  Opened valve one full turn at 2120 GMT.

6/20/13

CAM/JEM

Mentor requested the DMA blower assembly be removed from the Ambient State (AS) and swapped in for the failed downstream blower in the TDMA.  Removed both blower assemblies and installed AS blower assembly in TDMA. Capped off the open AS lines.  Powered on TDMA system and notified mentor it was ready to be started and checked for proper operation.  Mentor started TDMA and advised it appears to be working well.  Will need to purchase a new blower to replace the failed unit and then reinstall in AS.  TDMA RTS at 1645 GMT.

6/19/13

CAM

Instrument mentor requested assistance troubleshooting upstream DMA blower.   T/S found upstream flow meter was clogged with NaCl residue.  Removed flow meter and cleaned.  Replaced meter and upstream DMA blower RTS status O.K. at 2010 GMT.  Mentor called back later and indicated the downstream blower now has a problem.   Downstream blower box is extremely hot to touch.   T/S found no obstructions in downstream flow meter.  Suspect the downstream blower has failed.  Mentor will discuss the issue with colleagues and advise of next steps to take.  Shut TDMA down at 2100 GMT pending further instructions.

6/18/13

CAM

Mentor reported the counts were unusually high over the weekend and asked the HV switched be turned off to help determine if there is water in the lines.  Turned HV switches off at 2000 GMT.  Mentor called later and said the test did not appear to show moisture and requested HV turned back on.  Turned HV switches on at 2100 GMT.  Advised mentor that the landowner started cutting the wheat over the weekend, making a couple of passes near the AOS trailer.  Mentor suspects that could account for high counts during normal sampling, but counts were also high during calibration.  Mentor still looking into problem.

6/4/13

CAM

Noticed atomizer solution had been leaking all over the bottom of the TDMA enclosure as evident by the white NaCl residue left behind.  Mentor set up the midnight calibration to run during the day so we could watch and see where the leak was.  Solution return line was leaking around the bottle lid.  Removed lid for inspection and found the return line was recessed into the underside of the lid.  This allowed the returning solution to adhere to the underside of the lid and leak out around the bottle neck rather than dripping into the bottle.  Slid the Swagelok ferrule farther up the return line so the line would extend about 1/4” into the bottle.  Repeated the calibration and the returning atomizer solution now drips into the bottle as it should and no longer leaks.  RTS status O.K.

5/7/13

KT

The gauge for the intake line pump was reading 0. After inspection it was found that the intake line that connects to the pump outside the AOS building had separated from the pump. Re-installed hose line into pump. Power pump on, the gauge reading returned to normal. RTS at 1446 GMT.

4/18/13

KT

Instrument mentor requested that the new Proportional Flow Control Valve be installed. At 2000 GMT stopped computer program and powered down TDMA system. At 2005 GMT removed and replaced  Proportional Flow Control Valve. At 2040 GMT powered TDMA system on and started computer. RTS  at 2100 GMT. Old Valve SN# 10064, New Valve SN# 14195. Valves were made by Teknocraft INC.

4/15/13

KT

Instrument mentor requested that the RH hose be removed and replaced as well as Atomizer solution be added. At 2000 GMT removed and replaced the RH hose and also added Atomizer solution to the instrument. RTS at 2115 GMT.

4/9/13

KT

Instrument mentor requested that the Reservoir Heater be removed and replaced. At 1400 GMT stopped the program, removed and replaced the Reservoir Heater. At 1427 GMT started the program and gave mentor a call to notify him that the changes that were requested were made. RTS at 1430 GMT.

4/4/13

CW/PD/CM

Ambient state system reinstalled and all power, data and sample line connections completed.

3/21/13

KT

The TDMA computer program was not running. At 1311 GMT rebooted the computer. The computer came online and began normal operations. RTS at 1326 GMT.

3/15/13

PAD

Computer has power but is not running.  Disconnected serial cable from APS.  Pressed power button on computer but it would not power up.  Disconnected computer power cord and reconnected.  Powered up computer.  Reconnected APS serial cable.   At 1800 GMT, opened computer to verify cooling fans are running.  Blew out dust from inside computer case with compressed air.  RTS at 1427 GMT. 

3/14/13

PAD

Computer has power but is not running.  Disconnected serial cable from APS.  Pressed power button on computer but it would not power up.  Disconnected computer power cord and reconnected.  Computer powered up after cable was connected and program started correctly.  Reconnected APS serial cable.  RTS at 1548 GMT.

2/26/13

PAD

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Power was restored at 1920 GMT.  Filled reservoir with water and powered up reservoir heater.  Notified Mentor.  Started measurements on APS instrument.  Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/13/13

CAM/KT/PAD

Existing stairs and deck above AOS trailer are scheduled to be R/R with metal stairs and deck.  TDMA ambient state needs to be removed so existing stairs can be torn down.  Shut down TDMA at 1851 GMT.  Removed all tubing, BNC cables, power and comm cables running from TDMA to ambient state.  Capped vacuum line inside AOS trailer that ran to ambient state.  Placed HEPA filter on ambient state inlet at bulkhead of TDMA.  Started TDMA program at 1943 GMT.  Removed ambient state and took to shipping so unit could be cleaned and stored until re-install.  RTS at 2100 GMT.

2/7/13

CAM

Received a new air compressor for the TDMA.  Removed the lab compressor that had been temporarily installed on 1/28 and replaced with the new compressor.  RTS at 2220 GMT.

1/29/13

PAD/KT

CE44 and VE44 vacuum pumps have dropped to -14 inHg and -13 inHg respectively.  Pumps need to be changed out with rebuilt pumps.  Stopped program and R/R pumps.  Old: Thomas, model 2688 CE44, SN 010500000564.  model 2688 VE44, SN 060500002019.  New: Thomas, model CE44, SN 100500000844.  model VE44, SN 040900004569.  Vacuum pressures now show -16.5 inHg and -18.5 inHg.  RTS at 1600 GMT. 

1/28/13

CAM

The air compressor has failed again.  This is the second identical new compressor that has failed.  Removed failed compressor and temporarily installed one from the electronics lab.  Will work with mentor to monitor a calibration cycle and see how much air is required and order a suitable replacement.  RTS at 2100 GMT with compressor from lab.

12/24/12

PAD/JEM

24V power supply inside TDMA enclosure recently failed.  Temporary power supply from lab was installed and used until replacement was obtained.  Replacement power supply has been delivered to CF and needs to be installed.  Stopped program and shut down TDMA system.  Installed power supply and restarted system.  Sola Model SDN 10-24-100P, SN 1204789601.  RTS at 1508 GMT.

12/20/12

KT

The instrument power supply was not working. After trouble shooting, the conclusion was that the 24V power supply was bad and needed to be replaced. Installed temporary adjustable power supply until spare is available. RTS at 2000 GMT.

12/20/12

JEM

The UPS batteries need replaced.  Removed the battery pack from the UPS and replaced the batteries.  Tested and reinstalled the battery pack.

12/4/12

CAM

The air compressor has failed.  Motor runs continuously but does not compress air.  Returned compressor under warranty to point of purchase and exchanged for new one.  Installed new compressor and RTS at 1700 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  TDMA was gracefully shut down at 1312 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted TDMA and RTS at 2115 GMT.

11/16/12

KT

Work order requested that the air compressor be removed and replaced. Removed and replaced air compressor at 1520 GMT. Campbell Hausfeld #FP209402. RTS at 1555 GMT.

11/13/12

PAD/KT

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  At 2000 GMT, found system had restarted but RH up and down input values were showing -0.250.  Most other values shown were zero.  Disconnected APS that had caused similar readings after system restart with no change in status after rebooting computer.  T/S with mentor.  At 2250 GMT, verified indicator light on power supply for sheath blowers was on.  Turned switches on bulkhead on that were previously shut off during T/S and sheath blowers powered up.  Suspect power supply did not restart correctly after power was restored.  Powerstrip that supplies power supply was reset shortly before sheath blowers powered up.  RTS at 2315 GMT.

9/19/12

PAD

Tubing to CE pump was broken during installation of LABLE instrument cables into AOS building on 9/18/12, see CM report for more details.  After tubing was repaired, received word from mentor to restart instrument.  CF experienced power outage between 0528-1535 GMT.  At 1620 GMT started Labview program but all input values were zero except RH up and RH dn which were -0.250.  Cursor was moving erratically. Left voicemail with mentor who later returned call.  At 1745 GMT, stopped/ restarted program with no change.  Rebooted TDMA computer with no change.  Troubleshooting found APS instrument is interfering with Labview program starting correctly.  Stopped APS program and disconnected serial cable from APS to TDMA computer.  Labview program started correctly.  Reconnected serial cable and started measurements on APS.  RTS at 1800 GMT.

9/19/12

KT

The CE44 pump value was reading 0.  During inspection of the pumps located outside on the south end of the AOS building it was found that the tubing that runs from the pump to CE44 was broken. Connected the hose back to the pump but left system off and called instrument mentor to see if instrument mentor would rather start system up differently since pump had not been running. Waiting to hear from instrument mentor at this time.  RTS at 1450 GMT.

8/30/12

PAD

Repaired CPC has been returned from vendor and needs to be installed.  Butanol was drained from CPC currently in use yesterday and allowed to dry overnight. Stopped program at 2058 GMT.  R/R *CPC.  Started program and monitored system.  RTS at 2138 GMT.  Old TSI Model 376200, SN 70415196.  New TSI Model 3762, SN 36, WD 54676.

8/29/12

PAD

CPC has been returned from being repaired and needs to be installed.  Butanol needs to be drained so CPC currently in use can dry out before being removed. Drained butanol at 1939 GMT.  Instrument will be allowed to run 24 hours before removal.  RTS at 1954 GMT.

8/20/12

PAD

Ion static strips are due to be changed.  Shut system down at 1533 GMT.  R/R ion static strips.  Restarted program.  RTS at 1548 GMT.  Removed static strips have been taken to shipping dept to be returned to vendor for disposal. 

2/7/12

PAD

Mentor requests static strips replaced.  Strips were ordered, arrived, and need to be installed.  Stopped program at 2108 GMT.  R/R two NRD Staticmaster 1U200 ionizing units.  Reconnected tubing and restarted program.  Removed strips have been taken to shipping dept to send back to vendor for disposal.  RTS at 2128 GMT.

2/1/12

PAD

TDMA is used for calibration of the ACSM instrument.  Stopped program and changed atomizer fluid to that used for ACSM calibration.  Connected instruments with nonconductive tubing.  Worked with mentors of both instruments to complete calibration.  After calibration, shut down TDMA, changed atomizer solution back to NaCl solution,disconnected from ACSM instrument and performed weekly PM while instrument was down.  Restarted program.  RTS at 1615 GMT.

1/31/12

PAD/KT

Humidity sensor in TDMA ambient state is not reading correctly.  At 1935 GMT, it was showing -5% RH.  Replacement probe was ordered and mentor requests probe installed.  Shut down TDMA instrument at 1935 GMT.  Removed access panel from side of Ambient state and R/R humidity probes with attached circuit boards.  Installed access panel and cover dome.  After install, RH is still reading -2% RH.  Notified mentor.  Old Vaisala Model HMM22DCB, SN A2710052.  New Vaisala Model HMM100, SN G4820026.  RTS at 2030 GMT.

1/18/12

PAD/KT

Recently instrument has been unable to reach humidity set points.  Mentor suspects a leak in system and requests connections and tubing checked at humidifier and water catch container.  Pulled back insulation on tubing connecting water reservoir and "humidity in" connector on bulkhead.  Tubing is clear but badly discolored.  Found debris in tubing that had line mostly plugged where tubing connects to hose barb connector at bulkhead.  Removed tubing and cleaned connectors at bulkhead and reservoir cap.  Replaced tubing and restarted system.  Notified mentor.  RH values are now close to reaching setpoints.  RTS at 1549 GMT.

1/5/12

PAD

Mentor requests atomizer solution replaced with distilled water and 147 nm calibration spheres so a calibration can be performed to system.  Shut down system at 1401 GMT and dumped, flushed reservoir.  Added distilled water and 2 drops of sphere solution.  Installed reservoir and notified mentor via voicemail.  RTS at 1416 GMT.  At 2023 GMT, notified by mentor calibration was successful.  Emptied atomizer reservoir and filled with NaCl atomizer solution.  RTS at 2038 GMT.

12/14/11

PAD

Upstream DMA flowmeter was replaced on 12/7/11 and cleaned on 12/12/11.  Mentor request original flowmeter installed for testing to see if cleaning corrected flow issues.  Shut down system and R/R flowmeter at 1555 GMT.  Notified mentor.  Old Model M-11SLPM-D/5M, SN 48392.  New Model V-100LPM-D, SN 24412.  RTS at 1625 GMT.

12/12/11

PAD

Upstream DMA flowmeter, (Alicat Scientific Model V-100LPM-D, SN 24412), was removed on 12/7/11 and replaced with test flowmeter.  At that time, observed buildup of white powder inside plates of flowmeter.  Mentor requested electronics removed and plates cleaned inside flowmeter.  Removed electronics box and removed/cleaned plates with distilled water.  Dried plates with cloth and blew with canned air.  After reassembly, connected flowmeter to dry air source in Optical trailer to remove any residual moisture.  TDMA instrument was not affected during cleaning of flowmeter. 

12/7/11

PAD

Mentor requested voltage checked at BNC connectors and DMA flow controllers replaced.  Checked voltage with mentor via phone.  Shut down system and replaced test flow controller on downstream DMA with original.  Removed Model M-100SLPM-D/5M, SN 48392 and installed Model V-100LPM-D, SN 24414.  Removed upstream DMA flow controller and installed test controller.  Old Model V-100LPM-D, SN 24412, New Model M-100SLPM-D/5M, SN 48392.  Notified instrument mentor.  RTS at 2200 GMT.

12/5/11

PAD/KT

Mentor suspects issue with either downstream or upstream DMA flow controller and request voltage checked on BNC cable and if needed, replacement flow controller installed.  Checked voltage at analog input AI 22.  Voltage measured 4.8 and was increasing.  Mentor was on phone and advised to shut off Qsh_dn switch.  Voltage started to drop as soon as switch was shut off.  Mentor requested downstream flow controller replaced with spare provided.  R/R flow controller and notified mentor.  Mentor will run system through several scans and monitor data to see if downstream flow controller was in fact the problem.  Both controllers are Alicat-Scientific.  Old Model V-100LPM-D.  SN 24414.  New Model M-100SLPM-D/5m.  SN 48392.  RTS at 2007 GMT.

10/5/11

PAD

Remotely accessed computer and Labview program is not running.  Computer is slow to open and close windows or programs.  Rebooted computer.  Labview program started correctly.  RTS at 1324 GMT.

9/26/11

PAD

TDMA vacuum pumps need to be replaced with rebuilt pumps.  Shut down and R/R pumps at 2001 GMT.  Connected new pumps and restarted system.  Vacuum pressure has increased with rebuilt pumps to -18.5 inHg for CE and -17.5 inHg for VE pumps.  CE pump is a loaner from Mentor until ARM replacement pump can be rebuilt.  Currently waiting on parts.  Old CE44, Model 2688 CE44, SN 010500000564.  New SN 100500000844.  Old VE44, Model 2688 VE44, SN 0605000002019.  New SN 0409000004569.  RTS at 2016 GMT.

8/29/11

PAD

APS instrument was removed on 7/25/11 and sent for repairs.  Instrument needs to be RTS.  Installed APS instrument at 1634 GMT.  Connected power, comm, and inlet tubing.  Started instrument.  Consulting with mentor to verify measurement settings are correct.  RTS at 1649 GMT.  TSI Model 1109125, SN 70815064, WD 58069.

7/25/11

PAD

Instrument is not showing any values on the particle size distribution graph.  Total concentration is zero.  Internal filters were changed on 1/5/11 and instrument seemed to be operating correctly until 5/12/11 when no particles were shown on graph.  Troubleshooting including cleaning inner and outer orifices has failed to correct problem.  Now Aerosol flow is 1.23 l/min. and sheath flow is 3.74 l/min indicating filters need to be changed again.  After consulting with mentor, instrument needs to be sent to vendor to troubleshoot zero concentration and nothing shown on distribution graph.  Shut instrument down and removed at 1723 GMT.  RTS at 1738 GMT.  TSI Model 1109125, SN 70815064, WD 58069.

3/14/11

PAD

TDMA system was used today to do IE calibrations on ACSM instrument.  Atomizer solution was changed and instrument output was connected with ACSM.  Worked with TDMA and ACSM mentors during calibration.  Returned TDMA system to normal and restarted at 2043 GMT.

3/8/11

PAD

Notified by mentor TDMA system has not had successful nightly calibration since 3/2/11.  Troubleshooting with mentor indicated potential blockage of air or atomizer solution in aluminum block located above solution reservoir.  Removed aluminum block and found internal blockage where solution enters.  Unable to dislodge with canned air, but was able to remove with small pick and then clear with air.  Installed block and reconnected all lines.  RTS at 2030 GMT.

2/15/11

PAD

Circulating fan inside TDMA enclosure failed approximately one month ago.  Replacement was ordered, received, and needs to be installed.  Removed screen and power cord from old fan and installed on new.  Installed new fan in enclosure at 1535 GMT.

12/21/10

PAD

R/R vacuum pumps for TDMA system at 2034 GMT.  Model CE44, old SN 2591, new SN 0564.  Model VE44, old SN 4569, new SN 2019.  RTS at 2054 GMT.

8/18/10

PAD

Notified by mentor ambient state RH probe is showing negative values for humidity.  Mentor requests voltage checked at BNC connector from probe.  At 1437 GMT, shut down system and removed ambient state cover.  Voltage measured 4.58V DC at 1459 GMT when actual RH was 85%.  It measured 3.98V DC at 1517 GMT when actual was 71%.  Cleaned screen for enclosure and installed cover.  Notified mentor.  Started system at 1535 GMT.

7/6/10

PAD/CAM

Storms moved through CF area on 7/5/10. At 1333 GMT, observed TDMA UPS had failed.  All lights on UPS are lit except uneven load indicator light and first (bottom) light on battery level indicator bar.  UPS is not responsive to on/off buttons.  Shut off power on UPS with main power switch and indicator lights remain lit.  Disconnected internal batteries with no change in status when batteries were reconnected.  Notified instrument mentor.  Removed in rack UPS, model 2200, and installed two model 1000 UPS for temporary use until main UPS can be repaired.  Restored power and started TDMA system.  Notified mentor who can now remotely access instrument computer.  RTS at 1633 GMT.

3/11/10

PAD

Repaired upstream DMA blower has been returned from mentor and needs to be installed.  Installed blower at 2210 GMT and informed mentor.  Mentor remotely started instrument.  RTS at 2230 GMT.

3/8/10

PAD

At 2010 requested by mentor to TS upstream DMA blower not working.  Mentor remotely accessed system and reports blower indicates it is on but is not registering on the flowmeter.  TS found blower had failed.  Mentor later requested blower removed and sent for repairs.  At 2105 GMT, removed blower.  RTS at 2125 GMT.

2/11/10

PAD

Notified by site monitor TDMA system is not operating.  Troubleshooting found system has power, but program is not running.  Restarted program and notified instrument mentor.  RTS at 2329 GMT.

10/20/09

PAD

Notified by mentor TDMA system is not reaching humidity set points.  Found failed heater in humidifier water reservoir.  Removed 100W heater and installed spare 75W heater.  Heater is working and humidity should increase as water warms.  RTS at 1616 GMT.

10/19/09

PAD

Notified by mentor that system has not been showing counts on nightly calibration using atomizer solution.  Mentor request manual valve opened a couple of turns to remove any buildup on end of needle valve.  Stopped system and adjusted valve as instructed at 1950 GMT. 

10/19/09

PAD

Error displayed:  Error 7 occurred at New File.  Loop # for today is 0.  Stopped and restarted ground_tdma.vi program.  Program restarted correctly.  RTS at 1440 GMT.

10/1/09

PAD

At 1404 GMT, observed program was not running.  Program had been stopped on 9/30/09 for weekly PM and was not restarted.  Started program and notified mentor who lowered RH setpoint until instrument reaches an optimal operating temperature.  Program started correctly and no other issues are known at this time.  RTS at 1419 GMT.   

8/17/09

PAD

Labview error message shown:  Error 7 occurred at new file.  Program time stamp shows 17 Aug 00:08:09.  Loop # for today is 0.  Stopped and restarted ground_tdma.vi program.  Program restarted correctly with no errors shown.  RTS at 1813 GMT.

8/5/09

PAD

Labview error is displayed:  Error 7 occured at New File.  Continue or Stop.  Labview program is shown but not active behind error message.  Instrument status shows loop # = 20, loop # today = 0.  When I selected continue, a second error was displayed:  Error 7 occurred at Open/ Create/ Replace File in Write spreadsheet string.vi>Write to spreadsheet file (DBL).vi>ground_tdma.vi.  Selected Stop and cleared error.  Stopped program and performed weekly PM on instrument.  Started Labview program with no errors shown after restart.  Instrument status now shows loop # = 1.  Loop # today = 3.  RTS at 1612 GMT.

7/15/09

PAD

Mentor request vacuum pumps replaced so current pumps can be rebuilt.  Shut down system at 2050 GMT, plumbed new pumps and R/R CE44 and VE44 pumps. RTS at 2120 GMT.

3/12/09

PAD

Mentor request static ion strips replaced in TDMA system.  Shut system down at 1822 GMT.  R/R ion strips and cleaned strips housing.  Reassembled housing and started system.  Notified instrument mentor.  RTS at 1845 GMT.

1/28/09

PAD

VE44 pump shows 0 inHg of vacuum.  Suspect vacuum tubing is frozen between AOS trailer and pump shelter.  Blocked exhaust vent located close to tubing entering pump shelter and increased temperature setting on heater located in shelter.  At 1539, line started to thaw out and flow returned to normal.  RTS at 1540 GMT.

10/27/08

PAD

Mentor requests CE44 vacuum pump for TDMA system replaced.  CE44 pump currently shows -13 inHg vacuum.  Shut down and R/R pump at 1943 GMT.  After startup new pump shows -17 inHg.  RTS at 1958 GMT.

10/14/08

PAD

Notified by mentor on 10/14/08 that TDMA data was very unusual and he had shut instrument down.  Mentor requested instrument visually inspected for any causes.  Found CPC butanol reservoir empty.  Filled reservoir and contacted mentor.  It was determined that after butanol reservoir emptied, controlled fill valve will remain open and instrument will pull inlet air through reservoir giving sporadic false particle count readings.  After restart, instrument started showing particle counts that are in range.  RTS at 1535 GMT.

9/29/08

PAD

Notified by TDMA mentor CCN count is low.  Troubleshooting found failed vacuum pump that pulls air that supplies TDMA system, CCN, and APS instruments.  R/R pump and RTS at 2120 GMT.

8/19/08

PAD

TDMA VE44 vacuum line is showing -15.5 inHg.  Found temporary hose attached to VE pump at last pump installation  had partially collapsed.  Removed hose and installed correct fitting provided by mentor.  Restarted pump and line now shows -16.5 inHg vacuum.  RTS at 1603 GMT.

8/1/08

PAD

Mentor request TDMA vacuum pumps replaced with pumps provided.  Stopped system and removed pumps at 1540 GMT.  Could not remove inlet Swageloc connector needed for replacement pump without damaging threads.  Could not locate replacement fitting on or off site.  Used other fittings available as temporary fix and installed pumps.  Restarted system at 1820 GMT.

7/9/08

PAD

CF experienced power loss from 0745 to 1358 GMT.  Computer is not running.  Reset UPS for TDMA system.  Rebooted AOS computer, restarted CPD, and started humidifier.  RTS at 1445 GMT.

6/6/08

PAD

Ambient state HV is fluctuating .  Mentor request BNC cable ran from TDMA instrument enclosure out to Ambient State to test cable or short issue.  Connected cable as instructed for testing.  After testing, mentor requested BNC cable installed with other cables between Ambient State and TDMA enclosure.  Pulled cable through conduit and connected to HV input at Ambient State and AO2 connector on panel inside TDMA enclosure.  RTS at 2035 GMT.

5/27/08

PAD

Mentor requests Ambient State high voltage power supply replaced with spare provided.  At 1905 GMT, Mark Klassen and I removed cover and R/R high voltage power supply.  Notified instrument mentor.  RTS at 1925 GMT.

5/19/08

PAD

Voltage at ambient state DMA has been fluctuating.  Mentor requests voltages measured at HV power supply input.  At 1845 GMT, removed ambient state cover and accessed bnc connector going to HV power supply.  Measured voltages as instrument was cycling through ambient state mode and reported voltages to mentor.  Voltages at HV power supply input measured correctly.  RTS at 1945 GMT.

5/5/08

PAD

Notified by mentor Ambient State HV supply is not ramping up in voltage.  Mentor request HV supply troubleshoot.  At 1840 GMT, Mark Klassen and I removed cover to Ambient State and checked voltages at control cable connector as per mentor's instructions.  Voltages were correct.  Reseated HV cable between power supply and DMA.  At 1950 GMT, shut down instrument and checked voltages at BNC connector on Ambient State control box as per mentor's instructions.  Voltages were correct.  Mentor remotely shut down and restarted system.  RTS at 2030 GMT.

4/24/08

PAD

Mentor request new comm cable installed connecting AOS computer, TDMA computer, and CCN instrument.  Disconnected data cable at CCN instrument to AOS hub and installed cable provided by mentor.  Connected cable to CCN, AOS hub and TDMA hub.  CCN will now send data to AOS computer but AOS computer cannot transmit to CCN instrument.  Troubleshooting at this time.  RTS at 1650 GMT.

4/8/08

PAD

Power outage at CF between 0709 and 0910 GMT.  At 1420 GMT, reset UPS and computer and air compressor came on.  Program started after computer boot up.  Stopped program and notified instrument mentor so system could be operated without humidity.  Mentor remotely started instrument and is drying system out at this time.  RTS at 1440 GMT.

4/2/08

MK

The TDMA UPS is faulting with bad battery light on. Removed battery pack, replaced all eight 12 volt batteries and reinstalled battery pack, cleared error's. The UPS is back to normal operation.

1/10/08

PAD

Mentor is troubleshooting arcing in downstream DMA and request HV input connector removed and an O-ring installed on HV connector.  Shut system down at 1657 GMT and removed HV connector.  Took pictures of connector and removed teflon spacer and heat shrink.   Installed o-ring supplied by mentor and installed.  No improvement was found.  At 1932 GMT, removed HV connector and installed a second o-ring and restarted system.  Mentor reported no improvement.  At 2110 GMT, removed connector and installed third o-ring.  RTS at 2125 GMT.

1/7/08

PAD

Mentor is troubleshooting leak in downstream DMA and request temp/ RH sensor checked.   At 1718 GMT, removed downstream DMA Temp/ RH sensor and put teflon tape around ferule and reinstalled.  No change detected.  Mentor requested sensor removed and cap placed on opening for testing purposes.  This was done at 2040 GMT.  No change detected.  At 2215 GMT, removed cap and installed temp/ RH sensor.  Turned HV down switch to on position.  Troubleshooting at this time.  RTS at 2230 GMT.

11/30/07

PAD

Replacement cap for downstream DMA received at CF and needs to be installed.  Mark Klassen and I installed cap sent by mentor on downstream DMA.  Notified instrument mentor who remotely started instrument to test.  No problems with instrument known at this time.  RTS at 1908 GMT.

11/29/07

PAD

Computer and instrument have power but are not running.  No lights are on computer, no activity on hard drive, and it does not respond to power button or reset button.  Removed power cord from back of computer and restored power.  Computer booted up and labview program started.  RTS at 1645 GMT.

11/21/07

PAD

TDMA has been down this week due to HV input short in downstream DMA cap.  Mentor request line plugged and power restored to instrument so some scans can still be done with upstream DMA.  Capped line to downstream DMA and restored power to instrument.  Monitored instrument after mentor remotely started.  RTS at 1740 GMT.

11/20/07

PAD

Mentor request radon gas analyzer shut down for 24 hours for background check.  Shut off analyzer and inlet valve at 2020 GMT.  Notified mentor.  RTS at 2045 GMT.

11/19/07

PAD

Mentor request HV input connector removed from Ambient State and placed in downstream DMA in order to get that part of the instrument RTS.  Mark Klassen and I removed connector from Ambient State.  Upon trying to install HV connector into downstream DMA cap, found more problems caused by arcing within DMA cap.  Removed DMA cap and found carbon trails and what looks like corrosion inside cap.  Notified mentor.  Mentor advised sending DMA cap for inspection and/ or repair.  RTS at 2210 GMT.

11/15/07

PAD

T/S ongoing issue with mentor concerning downstream DMA HV.  Mentor request top cap of downstream DMA removed and sent for inspection and/ or repair.  Shut down system and tried to remove top cap.  Inner tubing of DMA is attached to cap.  Consulted with mentor for instructions.  After receiving instructions, removed HV input connector from downstream DMA cap.  Connector shows carbon arc trail from short.  Sent for inspection and repair.  RTS at 2005 GMT.

11/14/07

PAD

Downstream DMA HV is sporadic.  Mentor request T/S problem via phone.  Shut down instrument at 1519 GMT.  Switched input cables to HV power supply with same results.  Measured voltage on input cable to both HV power supplies during test.  Mentor is monitoring instrument at this time.  RTS at 1635 GMT.

11/9/07

PAD

HV is dropping out on downstream DMA.  Mentor request input and output cables swapped between upstream and downstream HV power supplies.  Stopped program and shut down power to instrument.  Swapped input and output cables on HV power supplies.  Started program.  RTS at 2116 GMT.

11/8/07

PAD

Requested by mentor to T/S suspected problem with HV power supply.  Shut down program and power to instrument.  Removed and inspected BNC cable from HV power supply to TDMA.  No problems found.  Attached cable and restarted system.  Monitored systems through scans and notified mentor of results.  Mentor requested and instrument was shut down at 2054 GMT.  Waiting on repairs at this time. 

10/30/07

PAD

Scheduler Vi screen shows 29 Oct 20:19:43.  Temp and RH values are not updating.  Notified instrument mentor who will remotely access instrument.  RTS at 1610 GMT.

10/4/07

PAD

At 1620 GMT, found instrument and computer with power but not running.  Contacted mentor to verify instrument had not been remotely shut down.  Computer would not turn on with power switch.  Unplugged/ plugged in computer and it started.  After instrument start-up, sheath blower would not spin down.  T/S with mentor and removed BNC cable and grounded to side of case to remove charge.  Instrument restarted correctly.  RTS at 1650 GMT.

10/3/07

PAD

CF experienced loss of power on 10/2/07.  Power has been restored and instrument needs RTS.  At 1430 GMT, turned on computer and instrument rack.  Started CPD program.  Notified instrument mentor.  RTS at 1445 GMT.

10/2/07

PAD

CF experienced loss of power at 1945 GMT.  Utilities company informed us outage could last for hours.  Management decided to shut instruments down to prevent problems at start up.  At 2009 GMT, found computer and instrument had already shut down.  Removed power cord supplying power to instrument and computer.  RTS at 2024 GMT.

9/12/07

PAD

Mentor request vacuum pumps for TDMA system R/R with pumps recently sent to CF.  At 1552 GMT, stopped TDMA program and shut down pumps.  R/R CE and VE pumps located in pump rack at AOS trailer.  After install, CE pump shows -19 inHg, VE pump shows -16.5 inHg.  Notified instrument mentor.  RTS at 1612 GMT.

8/30/07

PAD

CPC butanol reservoir is low.  Supply reservoir has butanol.  Contacted mentor and mentor remotely accessed computer.  Program was locked up.  Mentor remotely rebooted computer.  Program restarted correctly and CPC reservoir filled with butanol.  Removed panel from enclosure for 10 minutes to allow cooling of instruments.  RTS at 1930 GMT.

7/3/07

PAD

R/R water reservoir 75W heater with 100W heater sent by mentor.  RTS at 2055 GMT.

7/3/07

PAD

Mentor requests ambient state sheath blower R/R with repaired blower.  At 2015 GMT, stopped program and R/R blower in ambient state enclosure.  RTS at 2040 GMT.

6/5/07

PAD

Pump for line identified with red tape shows 0 in Hg.  This is the line that supplies TDMA inlet.  Removed pump at 1850 GMT and installed rebuild kit.  Pump now shows 17 in Hg.  RTS at 1940 GMT.

5/8/07

PAD

Notified by mentor atomizer is not working properly during daily calibration of TDMA instrument.  Mentor request aluminum mixing block located above atomizer reservoir removed and inspected for possible causes.  At 1926 GMT, stopped program and shut down instrument.  Removed atomizer mixing block and inspected.  Found orifice behind inlet connector was partially plugged and not letting enough airflow to draw atomizer solution into mixing block.  Cleaned orifice and tested system outside instrument enclosure.  Mixing block seems to be working correctly at this time.  Notified instrument mentor.  RTS at 2046 GMT.